Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

children not selected

Expert replies
by ruplun » Fri Jul 29, 2011 5:14 am
From a group of 5 men, 4 women and 3 children, if 4 persons are selected at random, what is the probability of no child being selected?
Join the discussion
Source: — Problem Solving |

by ColumbiaVC » Fri Jul 29, 2011 4:52 pm
Is it (9c4/12c4)=14/55 ?
Join the discussion

by tryingtobeat » Fri Jul 29, 2011 5:08 pm
or is it 41/55?
nm...

14/55 should be correct.
Join the discussion

by Bek » Fri Jul 29, 2011 7:26 pm
answer: 14/55

9/11*8/10*7/9*6/8=14/55
Join the discussion

by ruplun » Fri Jul 29, 2011 8:59 pm
the series is coming as 21/55 not 14/55...please explain
Join the discussion

by Bek » Fri Jul 29, 2011 9:18 pm
Sorry,

answer: 21/55

9/11 * 8/10 * 7/9 * 6/8 = 21/55

What is OA?
Join the discussion

by ruplun » Fri Jul 29, 2011 9:42 pm
sorry dont hav the correct ans wid me
Join the discussion

by Jim@Knewton » Sat Jul 30, 2011 12:16 pm
Good question Ruplun :!:
As with most other probability questions, there are a variety of ways to solve this, and here are a few:

1. The best and the easiest has been posted above by ColumbiaVC :!: (and Tryingtobeat :!:):
9c4/12c4 = 126/495 = [spoiler]14/55 = 0.254545[/spoiler]
where 9c4 = ways to select 4 from 9 (9 non-children, M or Wo does not really matter here)
...and 12c4 = all possible ways to select 4 from 12

2. Similarly, we can also do: (9/12)*(8/11)*(7/10)*(6/9) = 3024/11880 = [spoiler]14/55 = 0.254545[/spoiler]
This is also fairly intuitive, 9/12 ways to select the first 'non-child'
...so also 8/11 ways, 7/10 ways and 6/9 ways to select the 2nd, 3rd and 4th non-child(ren)

3. Long way - but this decomposes the problem analytically and is helpful for other variations of this and similar questions (such as pr(at least one child selected)):
Pr(No Child) = 1- pr(at least one child selected)
pr(at least one child selected)= Pr(3 Ch)+Pr(2 Ch)+ Pr(1 ch) [Ch= Child selected]
=>pr(at least one child selected)= (3*2*1*9*4c3 + 3*2*9*8*4c2 + 3*9*8*7*4c1)/(12*11*10*9)
Note that the denominator '(12*11*10*9)' has been factored out and 4cn represents the different ways in which 3, 2 or 1 child(ren) can be selected to fit the 4 spots.
=> pr(at least one child selected)= (3*2*1*9*4 + 3*2*9*8*6 + 3*9*8*7*4)/(12*11*10*9)
=> pr(at least one child selected)= [spoiler]8856/11880 = 41/55 = 0.74545[/spoiler]
=> Pr(No Child) = 1- pr(at least one child selected)= [spoiler]1- 41/55 = 14/55 = 0.254545[/spoiler]

4. Yes, even if the question were to find pr(at least one child selected), it would still be quicker to find Pr(all 4 non-children) and find pr(at least one child selected) by doing 1-Pr(all 4 non-children)

Hope this helps... :-) (Probability is always fun - that's where I myself make the most errors too!)
Best, Jim
Please "thank" this post if it helps you!
https://www.knewton.com/people/jims
Join the discussion

by Bek » Sat Jul 30, 2011 1:00 pm
Sorry, I miscalculated total number of people. It should 12, not 11.

answer: 14/55

9/12 * 8/11 * 7/10 * 6/9 = 14/55
Join the discussion